A machine can produce either nuts or bolts, but not both at the same time. The machine can be used at most 8 hours a day. Furthermore, at most 6 hours a day can be used for making nuts and at most 5 hours a day for making bolts. There is a $2 profit for each hour the machine makes nuts and a $3 profit for each hour the machine makes bots. How many hours per day should the machine make each item in order to maximize profit? What is the maximum profit? Please hurry!
Answer by Paul(988)

6(x+2)=5(x+3)
6x+12=5x+15
Simplify
x=3
6(3+2)=5(3+3)
6(5)=5(6)
30=30

Hence, the maximum profit should be $30.